A company xyz has a labelled mass of 500. It has a mean mass of 550 and standard deviation of 20. Assuming that masses are normally destributed at 550 at 5% significance level.


1
Expert's answer
2022-01-12T17:13:31-0500

"H_0:\mu=550" , mean mass is 550 kg

"H_0:\mu\neq550" , mean mass is not 550 kg


"z=\frac{\overline{x}-\mu}{\sigma}=\frac{500-550}{20}=-2.5"


critical value:

"z_{crit}=1.96"


Since "|z|>z_{crit}" we reject null hypothesis. Mean mass is not 550 kg